# DIFFA: Large Language Diffusion Models Can Listen and Understand

### v1

@@ 1 Introduction
Large Language Models (LLMs) have catalyzed a paradigm shift in artificial intelligence, advancing the frontiers of natural language understanding, computer vision, and multimodal reasoning. Extending these capabilities to the audio modality has given rise to Large audio-language models (LALMs), which integrate speech encoders with LLMs to enable spoken language understanding and generation. This direction holds transformative potential for applications such as voice assistants, human-computer interaction, and embodied AI.
Recent LALMs—such as Qwen2-Audio Chu et al. (2024), Qwen2.5-Omni Xu et al. (2025), SALMONN Tang et al. (2024a), and Audio-Flamingo Kong et al. (2024a)—predominantly adopt the autoregressive (AR) generation paradigm. These models rely on frozen LLM backbones (e.g., LLaMA 3 Dubey et al. (2024) or Qwen3 Yang et al. (2025)) paired with adapter modules that project acoustic representations into the LLM input space. Despite achieving strong performance, AR models remain limited by exposure bias, generation inefficiency, and restricted flexibility for bidirectional or partially conditioned inference.
To overcome these limitations, diffusion-based language models Austin et al. (2021); Shi et al. (2024) have emerged as a compelling alternative. By framing language generation as an iterative denoising process, these models enable non-autoregressive (NAR) decoding, parallel prediction, and greater controllability Shi et al. (2024). Pioneering work such as LLaDA Nie et al. (2025) has demonstrated that diffusion LLMs can achieve performance on par with AR models like LLaMA-3 Dubey et al. (2024), while offering improved robustness and training efficiency. More recently, LLaDA-V You et al. (2025) extends this approach to vision-language tasks, further confirming the generality and competitiveness of the diffusion modeling framework in multimodal settings.
@@ 2.1 Large Audio-Language Models (LALMs)
Large Audio-Language Models have emerged as powerful tools for processing spoken language in open-ended tasks. A key design pattern involves coupling a speech encoder (e.g., Whisper Radford et al. (2023)) with a pretrained LLM, often via lightweight adapters. For instance, Qwen-Audio Chu et al. (2023) and Qwen2-Audio Chu et al. (2024) integrate Whisper with Qwen via a unified input interface. SALMONN Tang et al. (2024a) proposes a dual-encoder structure and window-level Q-Former to better capture long audio contexts, while Audio-Flamingo2 Ghosh et al. (2025) achieves strong results by combining curriculum learning with dense alignment strategies. These models excel on benchmarks like MMAU and VoiceBench, yet rely heavily on autoregressive decoding.
Another direction focuses on general-purpose multimodal LLMs that treat audio as one modality among many. Early examples include PandaGPT Su et al. (2023) and AnyGPT Zhan et al. (2024), which convert diverse inputs (e.g., text, images, audio) into a common token space. Qwen-Omni Xu et al. (2025) builds upon this paradigm with a novel “Thinker-Talker” architecture and supports real-time spoken interaction across modalities.
Despite this progress, these approaches universally adopt autoregressive generation and have not explored non-AR architectures like diffusion models for audio-language understanding.
@@ 3.1 Preliminaries
$-\mathbb{E}_{t,p_{0},r_{0},r_{t}}\left[\frac{1}{t}\sum_{i=1}^{L^{\prime}}\textbf{1}[r_{t}^{i}=\textrm{M}]\log p_{\theta}(r_{0}^{i}|p_{0},r_{t})\right],$ (2)
@@ 5.1 Evaluation on MMSU
Table 5: Ablation study on model architecture and adapter design
@@
Table 6: Ablation study on the impact of different instruction data sources
@@ 5.2 Evaluation on MMAU
A breakdown by domain reveals that DIFFA performs best on speech-related tasks (59.46%), likely due to the ASR-focused pretraining on LibriSpeech. While models like Audio Flamingo 2 excel in music and sound, their speech performance lags behind (e.g., 44.74% for Audio Flamingo 2), highlighting the benefits of targeted speech-centric training.
@@ 5.3 Evaluation on VoiceBench
We evaluate DIFFA on VoiceBench, a comprehensive benchmark designed to assess the core capabilities of voice assistants across general knowledge, instruction following, and safety. The results are summarized in Table 4.
DIFFA achieves 29.57% on MMSU and 35.60% on OBQA, demonstrating emerging capabilities in spoken knowledge understanding. While this lags behind top models like GPT-4o-Audio (80.25%) and ASR+LLM (81.69%), it is notable given our model’s purely diffusion-based architecture and limited synthetic training data. These results suggest DIFFA can process factual spoken queries but still struggles with deeper semantic reasoning.
On IFEval, DIFFA scores 26.56%, outperforming models such as GLM-4-Voice (25.92%) and Qwen2-Audio (26.33%), and indicating a capacity for understanding simple spoken instructions. However, its performance remains far below systems like Kimi-Audio (61.10%) and GPT-4o-Audio (76.02%), reflecting a gap in handling complex audio-conditioned constraints.
In the safety-sensitive AdvBench, our model reaches 76.54%, competitive with many strong baselines. However, it still trails top-performing models like Kimi-Audio (100.00%) and GPT-4o-Audio (98.65%), underscoring the need for enhanced safety alignment during training.
VoiceBench results show that DIFFA establishes a promising baseline across key assistant tasks. Its performance confirms the feasibility of diffusion-based LLMs in spoken assistant settings, while highlighting future directions for improving semantic comprehension, instruction execution, and safety robustness.
@@ 5.4 Ablation Study
As shown in Table 5, replacing the autoregressive LLaMA 3.1 backbone with the diffusion-based LLaDA architecture leads to a substantial improvement across all metrics. Specifically, with dual adapters, the LLaDA variant achieves 51.71 on MMAU and 54.72 on MMSU, significantly outperforming its LLaMA counterpart. This highlights the advantage of masked diffusion language modeling for audio understanding tasks.
@@
We compare single and dual adapter configurations using the same Qwen3-based diffusion language model. The dual adapter setup—consisting of a semantic adapter and an acoustic adapter—consistently outperforms the single adapter baseline. Notably, dual adapters yield a +2.01 absolute gain on MMAU and a +3.16 gain on MMSU, demonstrating the benefit of jointly modeling both linguistic and acoustic features for audio understanding.
@@ 6 Conclusion
In this work, we introduce DIFFA, a diffusion-based large audio-language model, which leverages a frozen diffusion language backbone with lightweight dual adapters for audio understanding and instruction following. By adopting a two-stage training strategy and using purely synthetic instruction data, DIFFA achieves competitive performance on diverse audio-language benchmarks including MMSU, MMAU, and VoiceBench—despite relying on modest computational resources and no explicit supervised SFT data. As an initial exploration of diffusion-based language modeling in the audio domain, our results suggest that such models hold promise as an alternative to autoregressive LALMs, particularly for controllable and efficient generation. We hope this work opens new directions for building efficient, flexible, and controllable speech driven AI.

### v2

@@ 1 Introduction
Large language models (LLMs) have catalyzed a paradigm shift in artificial intelligence, pushing the frontiers of natural language understanding, computer vision, and multimodal reasoning Achiam et al. (2023). In the domain of speech and audio processing, large audio-language models (LALMs) have similarly benefited from these advances in LLMs. By bridging continuous acoustic signals with discrete linguistic representations, LALMs enable end-to-end modeling of spoken interaction Chu et al. (2023); Zhang et al. (2023). This capability not only advances fundamental research in speech understanding and generation, but also opens up practical opportunities for building more natural, robust, and versatile human-computer communication systems.
Existing LALMs typically follow two design paradigms. The first couples a speech encoder with an LLM, often through lightweight adapters that project continuous acoustic representations into the input space of the language model (e.g., Qwen2-Audio Chu et al. (2024), Audio-Flamingo Kong et al. (2024a)). The second discretizes audio into speech tokens via speech tokenizers and subsequently trains directly on these tokens under the LLM training paradigm (e.g., SpeechGPT Zhang et al. (2023), Moshi Défossez et al. (2024a)). Despite their strong results, both paradigms predominantly rely on autoregressive (AR) decoding, which suffers from well-known drawbacks such as exposure bias, slow generation, and limited flexibility for bidirectional or partially conditioned inference.
To address these limitations, diffusion-based language modelsAustin et al. (2021); Shi et al. (2024) have emerged as a promising alternative. By framing generation as an iterative denoising process, diffusion models support non-autoregressive decoding, parallel prediction, and improved controllabilityShi et al. (2024). Recent advances such as LLaDA Nie et al. (2025) demonstrate that diffusion LLMs can rival autoregressive counterparts like LLaMA-3 Dubey et al. (2024), while exhibiting stronger robustness and training efficiency. Furthermore, LLaDA-V You et al. (2025) extends this paradigm to vision–language tasks, confirming the competitiveness and generality of diffusion modeling in multimodal learning.
@@ 2.1 Large Audio-Language Models (LALMs)
Large audio–language models (LALMs) have recently emerged as powerful tools for spoken language understanding in open-ended tasks. Existing approaches generally fall into two paradigms. The first couples a speech encoder with a pretrained LLM, often via lightweight adapters to bridge acoustic and textual representations. Representative examples include Qwen-Audio Chu et al. (2023) and Qwen2-Audio Chu et al. (2024), which integrate Whisper through a unified interface; SALMONN Tang et al. (2024a), which introduces a dual-encoder and window-level Q-Former to better capture long audio contexts; and Audio-Flamingo2 Ghosh et al. (2025), which employs curriculum learning and dense alignment strategies. These systems achieve strong results on benchmarks such as MMAU and VoiceBench but remain tied to autoregressive decoding.
The second paradigm discretizes speech into tokens using quantizers or self-supervised encoders, and treats them as an additional input stream to the LLM (e.g., SpeechGPT Zhang et al. (2023), Moshi Défossez et al. (2024a)). While effective, this line of work also relies exclusively on autoregressive generation. To date, non-AR architectures such as diffusion models have not been explored for audio–language understanding.
@@ 3.1 Preliminaries
$\mathbb{E}_{t,p_{0},r_{0},r_{t}}\left[\frac{1}{t}\sum_{i=1}^{L^{\prime}}\textbf{1}[r_{t}^{i}=\textrm{M}]\log p_{\theta}(r_{0}^{i}|p_{0},r_{t})\right],$ (2)
@@ 5.1 Evaluation on MMSU
Table 5: Ablation study on model architecture and adapter design. All models use 8B Instruct version.
@@
Table 6: Ablation study on the impact of different instruction data sources. All models use 8B Instruct version.
@@ 5.2 Evaluation on MMAU
A domain-level breakdown shows that DIFFA achieves the highest performance on speech-related tasks (59.46%), likely benefiting from its speech-caption-focused training paradigm. In contrast, models such as Audio Flamingo 2 perform better on music and environmental sounds but underperform in speech understanding (e.g., 44.74% for speech), underscoring the advantage of targeted, speech-centric training.
@@ 5.3 Evaluation on VoiceBench
VoiceBench evaluates semantic understanding from audio-as-question prompts, covering knowledge, instructions, and safety—making it a rigorous benchmark for spoken query comprehension.
Despite being trained on only 960 hours of ASR data and 127 hours of synthetic instructions, DIFFA achieves 34.45% on SD-QA and 35.60% on OBQA, demonstrating promising capability in factual spoken QA. This is particularly notable when compared to models like Qwen2-Audio (35.72% SD-QA, 49.45% OBQA), which are trained on hundreds of thousands of hours of proprietary data. The results suggest that diffusion-based models, even with limited training, can capture core semantic structures in speech.
On IFEval, DIFFA reaches 26.56%, slightly surpassing Qwen2-Audio (26.33%) and GLM-4-Voice (25.92%), indicating a basic capacity for instruction comprehension from audio inputs. However, the performance gap with top models such as Kimi-Audio (61.10%) underscores the challenge of aligning audio-conditioned instruction execution without large-scale supervised tuning.
In AdvBench, DIFFA attains 76.54%, outperforming many strong baselines and approaching GLM-4-Voice (88.08). This highlights the potential of lightweight, diffusion-based models to learn safety-aligned behavior with minimal data.
In summary, DIFFA establishes a competitive baseline on VoiceBench despite using orders of magnitude less training data than competing systems. These findings validate the feasibility of diffusion-based LLMs for semantic audio understanding and offer a data-efficient alternative to current autoregressive paradigms.
@@ 5.4 Ablation Study
As shown in Table 5, replacing the autoregressive LLaMA 3.1 backbone with the diffusion-based LLaDA architecture leads to a substantial improvement across all metrics. Specifically, with dual adapters, the LLaDA variant achieves 51.71 on MMAU and 54.72 on MMSU, significantly outperforming its LLaMA counterpart. This highlights the advantage of diffusion language model for audio understanding tasks.
@@
We compare single and dual adapter configurations using the same LLaDA backbone. The single adapter setup uses only a semantic adapter aligned with the speech encoder’s output. The dual adapter adds an acoustic adapter that extracts low-level acoustic cues from intermediate encoder states. Results show that dual adapters yield a +2.01 gain on MMAU and +3.16 on MMSU, confirming the advantage of combining semantic and acoustic information for richer audio understanding.
@@ 6 Conclusion
In this work, we introduce DIFFA, a diffusion-based large audio-language model that combines a frozen diffusion language backbone with lightweight dual adapters for audio understanding and instruction following. Using a two-stage training strategy and fully synthetic instruction data, DIFFA achieves competitive performance on diverse benchmarks—including MMSU, MMAU, and VoiceBench—despite relying on only 960 hours of ASR data and 127 hours of synthetic instruction data, in contrast to models like Qwen2-Audio-Instruct trained on over 500K hours. As an initial exploration of diffusion-based modeling in the audio domain, our results suggest that such models offer a promising alternative to autoregressive LALMs. While our experiments are conducted on relatively small-scale open-source corpora, we plan to scale to broader and more diverse data in future work. We hope this study encourages further research into efficient, flexible, and controllable speech-driven AI systems.